# On the Hasse Principle for Conic Bundles over Even Degree Extensions

### v1

@@ 1. Introduction
Let $k$ be a number field and let $X\to\mathbb{P}^{1}_{k}$ be a conic bundle. Assume Schinzel’s hypothesis. Then, there exists an explicit finite set $S$ of quadratic extensions of $k$ such that for every quadratic extension $L/k\notin S$, then
@@
In 2009, Pete Clark coined the notion of a potential Hasse principle failure for smooth, geometrically irreducible varieties over number fields. In particular, given such a variety $V/k$, we say that it is a potential Hasse principle failure (PHPF) if there exists an extension $L/k$ such that $V$ fails the Hasse principle over $L$. Clark showed the existence of infinitely many curves which were PHPFs. It was also conjectured that for any positive genus curve $C/k$ with no $k$-rational points, $C$ is a PHPF, see [Cla05]*Theorem 1, Conjecture 4. 1.1 shows that most conic bundles with four geometric singular fibers are not PHPFs, although there do exist cases where new Brauer classes can give a Brauer-Manin obstruction over $L$. If the conditions of 1.1 are not met, there exists a quadratic extension $L/k$, where the singular fibers of $X_{L}\rightarrow\mathbb{P}_{L}^{1}$ lie over two closed points of degree $2$, over which the Hasse principle could possibly fail.
@@ 2.2. The Brauer group of a conic bundle
For any point $P\in\mathbb{A}_{k}^{1}$, let $\mathbf{k}(P)$ denote its residue field. The fiber $X_{P}$ degenerates to the union of two lines, $\ell_{P}$ and $\ell^{\prime}_{P}$, defined over $\mathbf{k}(P)(\sqrt{\alpha_{P}})$ for some $\alpha_{P}\in\mathbf{k}(P)^{*}$. Since $X$ is assumed to be minimal, if $P\in S$ is a degree $1$ point, then $\alpha_{P}$ cannot be a square in $k$, as both $\ell$ and $\ell^{\prime}$ would be (1)(-1)-curves defined over $k$. Note that the generic fiber of $X\to\mathbb{P}^{1}$ is a Severi-Brauer variety over $k(t)$ and therefore an element of $\operatorname{Br}(k(t))$, and $\alpha_{P}$ is the residue of this Brauer class at $P$. The group $\frac{\operatorname{Br}(X)}{\operatorname{Br}_{0}(X)}$ is then determined by $\mathbf{k}(P)$ and $\alpha_{P}$ for $P\in S$, as shown in the following lemma.
@@
Let $L/k$ be a finite extension such that the points of $S$ remain unchanged over $L$. For every $T\subseteq S$, let $\beta_{T}:=\prod\limits_{P\in T}\operatorname{Nm}_{\mathbf{k}(P)/k}(\alpha_{P})$. If $\beta_{T}\in k^{\times 2}\Leftrightarrow\beta_{T}\in L^{\times 2}$ for all $T\subseteq S$, then the map $\operatorname{Res}_{L/k}\colon\frac{\operatorname{Br}(X)}{\operatorname{Br}_{0}(X)}\to\frac{\operatorname{Br}(X_{L})}{\operatorname{Br}_{0}(X_{L})}$ is an isomorphism.
@@
where the last horizontal map sends $\varepsilon\in\mathbb{F}_{2}^{|S|}/(1,\dots,1)$ to $\prod\limits_{P\in S}\operatorname{Nm}_{\mathbf{k}(P)/k}(\alpha_{P})^{\varepsilon_{P}}$ and analogously for $L$, and the vertical maps are induced by $\operatorname{Res}_{L/k}$. In particular, the map $\mathbb{F}_{2}^{|S|}/(1,\dots,1)\to\mathbb{F}_{2}^{|S_{L}|}/(1,\dots,1)$ sends $e_{P}$, the basis vector corresponding to $P$, to the sum $e_{Q_{1}}+\cdots+e_{Q_{n}}$, where $P$ splits into $Q_{1},\dots,Q_{n}$ over $L$. The diagram above induces isomorphisms compatible with restriction
$\frac{\operatorname{Br}(X)}{\operatorname{Br}(k)}$$\ker\left(\mathbb{F}_{2}^{|S|}/(1,\dots,1)\to k^{\times}/k^{\times 2}\right)$$\frac{\operatorname{Br}(X_{L})}{\operatorname{Br}(L)}$$\ker\left(\mathbb{F}_{2}^{|S_{L}|}/(1,\dots,1)\to L^{\times}/L^{\times 2}\right)$}}
where the vertical maps are given by $\operatorname{Res}_{L/k}$. Since $S=S_{L}$ and using the explicit description of the map above, the map $\mathbb{F}_{2}^{|S|}/(1,\dots,1)\to\mathbb{F}_{2}^{|S_{L}|}/(1,\dots,1)$ is an isomorphism. The condition on $\beta_{T}$ precisely implies that the kernels will be equal, and therefore $\operatorname{Res}_{L/k}$ is an isomorphism as desired. ∎
@@ 3. Proofs of main theorems
Following this construction for each place $v\in\Omega_{k}$, we produce $(P_{w})\in\mathbb{A}_{L}$ satisfying

### v2

@@ 1. Introduction
Let $k$ be a number field and let $X\to\mathbb{P}^{1}_{k}$ be a conic bundle. Then, there exists an explicit finite set $S$ of quadratic extensions of $k$ such that for every quadratic extension $L/k\notin S$, we have
@@
In 2009, Pete Clark coined the notion of a potential Hasse principle failure for smooth, geometrically irreducible varieties over number fields. In particular, given such a variety $V/k$, we say that it is a potential Hasse principle failure (PHPF) if there exists an extension $L/k$ such that $V$ fails the Hasse principle over $L$. Clark showed the existence of infinitely many curves which were PHPFs. It was also conjectured that for any positive genus curve $C/k$ with no $k$-rational points, $C$ is a PHPF, see [Cla05]*Theorem 1, Conjecture 4. 1.1 shows that most conic bundles with four geometric singular fibers are not PHPFs, although there do exist cases where new Brauer classes can give a Brauer-Manin obstruction over $L$. If the conditions of 1.1 are not met, it is possible for there to exist a quadratic extension $L/k$, where the singular fibers of $X_{L}\rightarrow\mathbb{P}_{L}^{1}$ lie over two closed points of degree $2$, over which the Hasse principle could possibly fail.
@@ 2.2. The Brauer group of a conic bundle
For any point $P\in\mathbb{A}_{k}^{1}$, let $\mathbf{k}(P)$ denote its residue field. For $P\in S$, the fiber $X_{P}$ degenerates to the union of two lines, $\ell_{P}$ and $\ell^{\prime}_{P}$, defined over $\mathbf{k}(P)(\sqrt{\alpha_{P}})$ for some $\alpha_{P}\in\mathbf{k}(P)^{*}$. Since $X$ is assumed to be minimal, if $P\in S$ is a degree $1$ point, then $\alpha_{P}$ cannot be a square in $k$, as both $\ell$ and $\ell^{\prime}$ would be (1)(-1)-curves defined over $k$. Note that the generic fiber of $X\to\mathbb{P}^{1}$ is a Severi-Brauer variety over $k(t)$ and therefore an element of $\operatorname{Br}(k(t))$, and $\alpha_{P}$ is the residue of this Brauer class at $P$. The group $\frac{\operatorname{Br}(X)}{\operatorname{Br}_{0}(X)}$ is then determined by $\mathbf{k}(P)$ and $\alpha_{P}$ for $P\in S$, as shown in the following lemma.
@@
Let $L/k$ be a finite extension such that $S_{L}=S_{k}$. For every $T\subseteq S$, let $\beta_{T}:=\prod\limits_{P\in T}\operatorname{Nm}_{\mathbf{k}(P)/k}(\alpha_{P})$. If $\beta_{T}\in k^{\times 2}\Leftrightarrow\beta_{T}\in L^{\times 2}$ for all $T\subseteq S$, then the map $\operatorname{Res}_{L/k}\colon\frac{\operatorname{Br}(X)}{\operatorname{Br}_{0}(X)}\to\frac{\operatorname{Br}(X_{L})}{\operatorname{Br}_{0}(X_{L})}$ is an isomorphism.
@@
where the last horizontal map sends $\varepsilon\in\mathbb{F}_{2}^{|S_{k}|}/(1,\dots,1)$ to $\prod\limits_{P\in S}\operatorname{Nm}_{\mathbf{k}(P)/k}(\alpha_{P})^{\varepsilon_{P}}$ and analogously for $L$, and the vertical maps are induced by $\operatorname{Res}_{L/k}$. In particular, we have that the map $\mathbb{F}_{2}^{|S_{k}|}/(1,\dots,1)\to\mathbb{F}_{2}^{|S_{L}|}/(1,\dots,1)$ sends $e_{P}$, the basis vector corresponding to $P$, to the sum $e_{Q_{1}}+\cdots+e_{Q_{n}}$, where $P$ splits into $Q_{1},\dots,Q_{n}$ over $L$. The diagram above induces isomorphisms compatible with restriction
$\frac{\operatorname{Br}(X)}{\operatorname{Br}(k)}$$\ker\left(\mathbb{F}_{2}^{|S_{k}|}/(1,\dots,1)\to k^{\times}/k^{\times 2}\right)$$\frac{\operatorname{Br}(X_{L})}{\operatorname{Br}(L)}$$\ker\left(\mathbb{F}_{2}^{|S_{L}|}/(1,\dots,1)\to L^{\times}/L^{\times 2}\right)$}}
where the vertical maps are given by $\operatorname{Res}_{L/k}$. Since $S=S_{L}$ and using the explicit description of the map above, the map $\mathbb{F}_{2}^{|S_{k}|}/(1,\dots,1)\to\mathbb{F}_{2}^{|S_{L}|}/(1,\dots,1)$ is an isomorphism. The condition on $\beta_{T}$ precisely implies that the kernels will be equal, and therefore $\operatorname{Res}_{L/k}$ is an isomorphism as desired. ∎
@@ 3. Proofs of main theorems
Following this construction for each place $v\in\Omega_{k}$, we produce $(P_{w})\in X(\mathbb{A}_{L})$ satisfying
@@ 4. An additional condition on Châtelet surfaces
Declarations
Conflict of Interest Statement
The authors are not aware of any financial or non-financial interests related to this publication.
Data Availability Statement
No data was generated or analyzed for this paper, so data sharing is not applicable.